roll a fair dice twice let A be the event that the sum of the two roles equal 6 and B be the event that the same number comes up twice what is P(A/B)?

Grade:12

1 Answers

Khimraj
5 years ago
Possible outcome
(1,1)(1,2)(1,3)(1,4)(1,5)(1,6)(2,1)(2,2)(2,3)(2,4)(2,5)(2,6)(3,1)(3,2)(3,3)(3,4)(3,5)(3,6)(4,1)(4,2)(4,3)(4,4)(4,5)(4,6)(5,1)(5,2)(5,3)(5,4)(5,5)(5,6)(6,1)(6,2)(6,3)(6,4)(6,5)(6,6)
P(A) = 5/36
P(B) = 6/36
P(A and B) = 1/36
P(A/B) = P(A and B)/P(B) = 1/6